OpenStudy (anonymous):

Solve the inequality 2x^4 -3x^3-7x^2+12x-4 is greater than or equal to 0. give answer in interval notation.

OpenStudy (anonymous):

the factors are -1, 1/2, 1, 2 can you solve now?

OpenStudy (anonymous):

I don't know what that means. Could you expand?

OpenStudy (anonymous):

( x - 1 )( x - 1/2 )( x + 1)( x - 2 ) are the factors, can you solve further?

OpenStudy (anonymous):

I can graph it, I don't know how to write the answer in interval notation though

OpenStudy (anonymous):

\[ x \in ( -\infty, -1) U (\frac{ 1 }{ 2 },1) U(2 , \infty )\]

OpenStudy (anonymous):

maybe \[( x - 1 )(2 x - 1 )( x + 1)( x - 2 )\]

OpenStudy (anonymous):

 here all the brackets should be [] not () ;)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

;)